OpenStudy (lgbasallote):

Just checking my answer... Find the volume of the solid obtained by rotating the region bounded by \(y = \frac{1}{4} x^2\) and \(y = 5 - x^2\) about the x-axis. I got \(\frac{176\pi}{3}\)

OpenStudy (anonymous):

i got something else again...

OpenStudy (lgbasallote):

youuuu =____=

OpenStudy (anonymous):

it's sorta close to yours....

OpenStudy (anonymous):

wanna know what I got?

OpenStudy (lgbasallote):

this is washer method right?

OpenStudy (anonymous):

yep...

OpenStudy (lgbasallote):

then whyyyy =_=

OpenStudy (lgbasallote):

uhh phooey i forgot to integrate x^4/16 =_=

OpenStudy (lgbasallote):

\[\frac{868\pi}{13}\]

OpenStudy (anonymous):

umm...

OpenStudy (anonymous):

i hope you don't get mad at me.... your first answer is correct...

OpenStudy (anonymous):

yep.. you're mad.... cya...

OpenStudy (lgbasallote):

lol :PPPP

OpenStudy (lgbasallote):

seems i combine x^4/16 already haha

OpenStudy (anonymous):

i was squaring like this (g-f)^2 and not g^2 - f^2 ... imagine that... :P

OpenStudy (lgbasallote):

lol =))))

OpenStudy (anonymous):

now you are a master....